# Rethinking Cross-Attention for Infrared and Visible Image Fusion

### v1

@@ III-A Framework Overview
The goal of IV-IF is to obtain a comprehensive image that contains salient targets while preserving rich textural details. Thus, how to fully exploit discrepancy and common information present in the source images is a key determinant of the fusion performance. Motivated by the argument that the cross-attention mechanism is effective in extracting common features among images, we propose DIIM and ACIIM.
@@
To effectively obtain the discrepancy features between infrared and visible image features produced by the previous stage, we use a newly built cross-attention architecture, as illustrated in Fig. 2. It is given the $F_{ir}^{token}$ and $F_{vi}^{token}$ as input and the discrepancy information features as output. Specifically, to explore the long-range relationship of IV features, we partition the $F_{ir}^{token}$ and $F_{vi}^{token}$ into $s$ local feature segments as:
$Q_{1},\dots,Q_{s}=Partition(F_{vi}^{token}),$ (6)
$K_{1},\dots,K_{s}=Partition(F_{ir}^{token}),$
$V_{1},\dots,V_{s}=Partition(F_{ir}^{token}),$
@@
where $F_{vi}^{token}$ and $F_{ir}^{token}\in\mathbb{R}^{h\times w\times c}$, and $s=h\times w$. Afterward, we employ the linear layer to transform the token segments into the query Q, key K, and value V, which are three fundamental elements of the base Transformer. The linear projection can be expressed as:
$Q_{i}=Linear_{Q}(Q_{i}),$ (7)
$K_{i}=Linear_{K}(K_{i}),$
$V_{i}=Linear_{V}(V_{i}),$
where $i=1,\dots,s$, the $Linear(\cdot)$ is a linear projection operator that is shared across different segments.
@@
To explore the common information of infrared and visible image features with the consideration of long-term relationships, we use the dot-production attention layer to compute the similarity matrix between $Q_{i}$ and $K_{j}$ (i and j belong to 1 through s), and then multiply by the $V$ to infer the common information between Q and V. This process can be expressed as:
$CM_{QV}=softmax(\frac{Q_{1,\dots,s}K^{T}_{1,\dots,s}}{\sqrt{d_{k}}})V,$ (8)
@@
where $d_{k}$ is a scaling factor, which can alleviate the softmax function from converging to regions of minimal gradients as the dot product increases. Subsequently, we can easily obtain the discrepancy information between Q and V by removing the common information. This process can be represented as:
$DIM_{QV}=Linear(V-CM_{QV}).$ (9)
@@
To obtain complementary information from the IV images, we inject the discrepancy information into Q, which can be formulated as:
$F_{add}$ $=DIM_{QV}+Q$ (10)
$F_{diim}$ $=MLP(LN(F_{add}))+F_{add},$
where $LN(\cdot)$ represents layer normalization, $MLP(\cdot)$ represents multi-layer perception and $F_{diim}$ is the output of the DIIM.
@@
Figure 4: The architecture of alternate common information injection module (ACIIM).
As depicted in Fig. 3, although the obtained feature maps generated by the DIIM almost contain all the salient edge information, the background detail information is not completely preserved in the fusion features. Thus, it is essential to further enhance the fused features by incorporating common information from both the IV images. Following the DIIM, the proposed ATFuse alternately extracts common information reflecting the background details from the IV images through the ACIIM module. The structure of ACIIM is illustrated in Fig. 4. We first inject common information from visible images to enrich the fused information. Specifically, given that the segments of $F_{diim}$ are used as $Q_{1,\dots,s}$ and the segments of $F_{vi}^{token}$ are used as $K_{i,\dots,s}$ and $V_{i,\dots,s}$, the common information between $F_{diim}$ and $F_{vi}^{token}$ can be expressed as:
$CM_{vi}^{token}=softmax(\frac{Q_{1,\dots,s}K^{T}_{1,\dots,s}}{\sqrt{d_{k}}})V.$ (11)
Then, the common information $CM_{vi}^{token}$ between $F_{diim}$ and $F_{vi}^{token}$ is added with $F_{diim}$. The process can be formulated as:
$F_{add}$ $=Linear(CM_{vi}^{token})+Q,$ (12)
$F_{aciim}^{vi}$ $=MLP(LN(F_{add}))+F_{add},$
where $F_{aciim}^{vi}$ represents the output of the first ACIIM.
Afterward, the common information between $F_{aciim}^{vi}$ and $F_{ir}^{token}$ should also be injected into the fused features to enrich the fused features. This process follows the same manner as above Eq. 12.
III-C Loss Function
Since the proposed ATFuse is trained in an unsupervised end-to-end manner, the choice of loss function greatly affects the fusion performance. Considering the different imaging mechanisms of IV images, the loss function of the proposed method makes sure that the fusion results can preserve sufficient details and salient information. The loss function of our ATFuse can be defined as:
@@ III-C Loss Function
where ${L}_{pixel}$ and ${L}_{texture}$ represent the pixel loss and texture loss, respectively. $\gamma$ is a hyperparameter to balance these two loss terms.
@@
Motivated by [65], the texture detail of an image can be represented by maximum aggregation around the gradients. Thus, the texture loss is designed to regulate the gradients in the fused image as follows:
$\mathcal{L}_{texture}=\frac{1}{HW}\|\nabla I_{f}-max\{\nabla I_{ir},\nabla I_{vi}\}\|_{1},$ (14)
@@
where $\nabla$ represents the Sobel operator, which is used to calculate the gradient. $\|\cdot\|_{1}$ represents the $l_{1}$-norm. H and W denote height and width respectively. $max\{,\}$ represents the element-wise maximum selection.
@@
It is known that maximum value selection-based pixel loss weakens the importance of pixels in one of the source images, while the average pixel loss may reduce the saliency of fusion results. Therefore, to achieve a better compromise between the preservation of important information and the enhancement of salient information, we adopt a segmented pixel loss function to train the proposed framework. In this work, the importance of a pixel is defined as the product of the pixel value and its gradient value as
$pi=\nabla I\cdot I,$ (15)
@@
where $I$ represents the pixel value, while $\nabla I$ denotes the corresponding gradient value.
@@
The importance of each pixel in the two source images is assessed and categorized into two segments as follows,
$p_{part1}=\{p|(pi\geq pi_{ir}^{\alpha}\quad or\quad pi\geq pi_{vi}^{\alpha})\},$ (16)
$p_{part2}=\{p|(pi<pi_{ir}^{\alpha}\quad and\quad pi<pi_{vi}^{\alpha})\},$
@@
where $pi_{vi}^{a}$ and $pi_{ir}^{a}$ represent the top $\alpha\%$ important pixels. $p_{part1}$ denotes the first segment that consists of the most crucial pixels covering the top $\alpha\%$ important pixels from each source image. The remaining pixels are represented as the second segment $p_{part2}$.
@@
As depicted above, we apply different pixel loss functions to each segment to meet the tradeoff among pixels of different importance. The maximum value selection-based pixel loss is employed in $p_{part1}$, which can emphasize the saliency of these pixels in the fused image. Meanwhile, for $p_{part2}$, we use the pixel average loss to force the fused image to approximate each source image. The segmented pixel loss can be formulated as follows:
$\mathcal{L}_{part1}=\frac{1}{HW}\|I_{f}-max(I_{ir},I_{vi})\|_{1},\quad p\in p_{part1},$ (17)
$\mathcal{L}_{part2}=\frac{1}{2HW}(\|I_{f}-I_{ir}\|_{1}+\|I_{f}-I_{vi}\|_{1}),\quad p\in p_{part2},$
where $I_{ir},I_{vi}$ are source images. $I_{f}$ is the fused image. The final pixel loss function can be rewritten as follows.
$\mathcal{L}_{pixel}=\mathcal{L}_{part1}+\mathcal{L}_{part2},$ (18)
IV EXPERIMENTS
@@
In this section, we first introduce the datasets and implementation settings. Then, the comparative experimental methods and objective evaluation metrics are described. After that, the experiment results and discussion are reported. Furthermore, we conduct the ablation experiments and present the analysis of some hyperparameters and computational efficiency.
IV-A Datasets and Implementation Settings
In this study, we utilize the recent publicity datasets RoadScene [24] and MSRS [65], in which we select 201 pairs of IV images from each dataset. Then, these images are randomly cropped into sample patches with the size of $128\times 128$ for training. To train the proposed ATFuse, we utilize the AdamW optimizer to update the network parameters with an initial learning rate of $2\times 10^{-3}$, which is decreased by 50% each time after 50, 100, 200, and 400 epochs. The batch size is set to 16. The parameters $\alpha$ and $\gamma$ are set to 20 and 1.0, respectively. The configurations for all experiments are computed with an NVIDIA GeForce RTX 4090 GPU and 64 GB RAM. The framework is programmed on PyTorch.
IV-B Comparative Methods and Objective Evaluation Metrics
@@
Comparative Methods: To assess the superiority of our method, we select six state-of-the-art image fusion methods for comparison, including two Transformer-based methods: DATFuse[37] and SwinFusion[39], two autoencoder-based methods: DenseFuse[17] and RFN-Nest [18], an end-to-end CNN-based method: U2Fusion[24], and a GAN-based method: FusionGAN[25]. The source codes of all compared methods are publicly available and all parameters are configured as suggested in the original papers.
@@
Objective Evaluation Metrics: Six objective evaluation metrics are chosen to quantitatively evaluate the performance of all methods: Average gradient ($AG$), Entropy ($EN$)[66], Standard deviation ($SD$)[67], Spatial frequency ($SF$)[68], Visual information fidelity ($VIFF$)[69] and $Q_{abf}$. $AG$ measures the sharpness or clarity of an image. $EN$ is an objective measure of the amount of information contained in an image. $SD$ is a statistical theory-based standard deviation that reflects the degree of change in pixel brightness. $SF$ is a measure based on gradient distribution, which represents the rate of change of the grayscale of the fused image. $VIFF$ measures the quality of the fusion image according to visual information fidelity. $Q_{abf}$ is a non-reference quality evaluation index by estimating the degree of salient information from source images represented in the fused image. Higher values of these metrics correspond to better quality of the fused image.
Figure 5: Qualitative comparison of ATFuse with six state-of-the-art visible and infrared image fusion methods on the RoadScence dataset. (a) infrared image, (b) visible image, (c) DATFuse[37], (d) DenseFuse[17], (e) FusionGAN[25], (f) RFN-Nest[18], (g) SwinFusion[39], (h) U2Fusion[24], (i) Ours.
Figure 6: Using six metrics to quantitatively compare ATFuse with six existing fusion methods on the RoadScence dataset (twenty image pairs). The average of all metrics is shown in the legend.
IV-C Results and Discussion
1) Results on RoadScence Dataset: Fig. 5 shows five pairs of source infrared and visible images from the RoadScence dataset along with fusion results obtained through seven methods. Two localized areas in each image are magnified for better comparison. It is apparent that all seven methods can perform relatively good fusion effects. However, the other six methods still have some unsatisfactory performance. Obviously, results generated by DATFuse, SwinFusion, and ATFuse methods have more detailed information and salient information than those without using the Transformer architectures, since the attention mechanism of the Transformer can capture global information. Nevertheless, DATFuse and SwinFusion still have deficiencies when dealing with specific local information. As can be seen, the salient information of fusion results by DATFuse is relatively worse since it fails to extract distinctive information from the infrared image (see the red box in the second row and the green box in the first row). The SwinFusion method shows poor performance when processing low-saliency but information-rich regions (see the red box in the fourth row and the green box in the fifth row). One possible reason is that the pixel loss of SwinFusion makes the pixels of the fusion result completely tend to the maximum value of the source image, ignoring the information of the lower pixel values. Besides, DenseFuse, FusionGAN, RFN-Nest, and U2Fusion fail to preserve significant information from the infrared image well in the fusion results (see the green box in the first row and the red box in the second row). By contrast, it can be seen that our ATFuse maintains the salient information in all results, this is because the proposed ATFuse employs DIIM and ACIIM modules, which can generate results containing various types of information. Furthermore, the segmented pixel loss function enables our method to balance the retention of detailed information and salient information better.
Fig. 6 further reports quantitative results on the twenty image pairs of the RoadScence dataset by using six objective evaluation metrics. Results indicate that our proposed method performs the best average values on the four evaluation metrics (i.e., $AG$, $SF$, $VIFF$, and $Q_{abf}$). $EN$ and $SD$ rank second place as close to the best performance. Overall, our method has achieved satisfactory objective performance, which demonstrates the crucial role of the proposed network structure and the segmented pixel loss function.
Figure 7: Qualitative comparison of ATFuse with six state-of-the-art visible and infrared image fusion methods on the MSRS dataset. (a) infrared image, (b) visible image, (c) DATFuse[37], (d) DenseFuse[17], (e) FusionGAN[25], (f) RFN-Nest[18], (g) SwinFusion[39], (h) U2Fusion[24], (i) Ours.
Figure 8: Qualitative comparison of ATFuse with six state-of-the-art visible and infrared image fusion methods on the TNO dataset. (a) infrared image, (b) visible image, (c) DATFuse[37], (d) DenseFuse[17], (e) FusionGAN[25], (f) RFN-Nest[18], (g) SwinFusion[39], (h) U2Fusion[24], (i) Ours.

### v2

@@ 3.2 DIIM and ACIIM
The goal of VIF is to obtain a comprehensive image that contains salient targets while preserving rich textural details. Thus, how to fully exploit discrepancy and common information present in the source images is a key determinant of the fusion performance. Motivated by the concept that the cross-attention mechanism is effective in exchanging information between two different patches [52], we propose DIIM and ACIIM to fuse discrepancy and common features among the source images.
Figure 2: The architecture of the discrepancy information injection module (DIIM).
@@
To effectively obtain the discrepancy between the infrared and visible image features produced by the previous stage, we use a newly built cross-attention architecture, as illustrated in Fig. 2. It is given the $F_{ir}^{t}$ and $F_{vi}^{t}$ as input and the discrepancy information features as output. The $FN$ module consists of two operation steps. First, to explore the long-range relationship of IV features, we partition, through a partition operator $\mathcal{P}$, the $F_{ir}^{t}$ and $F_{vi}^{t}$ into $s$ local feature segments as:
$Q_{1},\dots,Q_{s}=\mathcal{P}(F_{vi}^{t}),$ (9)
$K_{1},\dots,K_{s}=\mathcal{P}(F_{ir}^{t}),$
$V_{1},\dots,V_{s}=\mathcal{P}(F_{ir}^{t}),$
@@
where $F_{vi}^{t}\in\mathbb{R}^{h\times w\times c}$, $F_{ir}^{t}\in\mathbb{R}^{h\times w\times c}$, and $s=h\times w$. Afterward, we employ the linear layer to transform the token segments into the query vectors ($Q$), key vectors ($K$), and value vectors ($V$), which are three fundamental elements of the base Transformer. The linear projection can be expressed as:
$Q_{i}=\text{Linear}_{Q}(Q_{i}),$ (10)
$K_{i}=\text{Linear}_{K}(K_{i}),$
$V_{i}=\text{Linear}_{V}(V_{i}),$
where $i=1,\dots,s$ and $\text{Linear}_{\cdot}(\cdot)$ are linear projection operators.
@@
To explore the common information $CM_{inf}$ of infrared and visible image features with the consideration of long-term relationships, we use the dot-production attention layer to compute the similarity matrix between $Q_{i}$ and $K_{j}$ ($i$ and $j$ from 1 to $s$), and then multiply by the vector $V$ to infer the common information. This process can be expressed as:
$CM_{inf}=\text{SoftMax}\left(\frac{Q_{1,\dots,s}K^{T}_{1,\dots,s}}{\sqrt{d_{k}}}\right)V,$ (11)
@@
where $\cdot^{T}$ is the transpose operator and $d_{k}$ is a scaling factor, which can alleviate the softmax function, $\text{SoftMax}\left(\cdot\right)$, from converging to regions of minimal gradients as the dot product increases. Subsequently, we can easily obtain the discrepancy information, $DIM_{inf}$, between the infrared and visible images by removing the common information. This process can be represented as:
$DIM_{inf}=\text{Linear}(V-CM_{inf}).$ (12)
@@
To obtain complementary information, $F_{add}$, from the IV images, we inject the discrepancy information into $Q$, which can be formulated as:
$F_{add}=DIM_{inf}+Q.$ (13)
Then, the final fused discrepancy features are obtained by layer normalization, $LN(\cdot)$, nonlinear mapping of the $F_{add}$ features, and skip connection operation:
$Z_{1}=MLP(LN(F_{add}))+F_{add},$ (14)
where $MLP(\cdot)$ includes two $\text{Linear}(\cdot)$ and one activation function $GELU(\cdot)$, and $Z_{1}$ represents the output of the DIIM.
@@
Figure 4: The architecture of the alternate common information injection module (ACIIM).
As depicted in Fig. 3, although the obtained feature maps generated by the DIIM almost contain all salient edge information, the background detail information is not completely preserved in the fusion features. Thus, it is essential to further enhance the fused features by incorporating common information from both IV images. Following the DIIM, the proposed ATFusion alternately extracts common information reflecting the background details from the IV images through the ACIIM module. The structure of the ACIIM is illustrated in Fig. 4. We expand Eq. (5). First, given that the segments of $Z_{1}$ are used as $Q_{1,\dots,s}$ and the segments of $F_{vi}^{t}$ are used as $K_{i,\dots,s}$ and $V$, the common information, $CM_{vi}^{t}$, between $Z_{1}$ and $F_{vi}^{t}$ can be expressed as:
$CM_{vi}^{t}=\text{SoftMax}\left(\frac{Q_{1,\dots,s}K^{T}_{1,\dots,s}}{\sqrt{d_{k}}}\right)V.$ (15)
Then, $CM_{vi}^{t}$ is added to $Z_{1}$ to compensate for the details of the fused discrepancy features. The process can be formulated as:
$F_{add}$ $=\text{Linear}(CM_{vi}^{t})+Q,$ (16)
$Z_{2}$ $=MLP(LN(F_{add}))+F_{add},$
where $Z_{2}$ represents the output of the first ACIIM.
Similarly, we compute the common information between $Z_{2}$ and $F_{ir}^{t}$ to extract the detail information from the infrared image according to Eqs. (15) and (16).
3.3 Loss Function
Since the proposed ATFusion is trained in an unsupervised end-to-end manner, the choice of a loss function greatly affects the fusion performance. Considering the different imaging mechanisms of IV images, the loss function of the proposed method must ensure that the fusion results can preserve sufficient details and salient information. The loss function of our ATFusion can be defined as:
@@ 3.3 Loss Function
where $\mathcal{L}_{pixel}$ and $\mathcal{L}_{texture}$ represent the pixel loss and texture loss, respectively. $\gamma$ is a hyperparameter to balance these two loss terms.
@@
Motivated by [53], the texture details of an image can be represented by maximum aggregation around the gradients. Thus, the texture loss is designed to regulate the gradients in the fused image as follows:
$\mathcal{L}_{texture}=\frac{1}{H\times W}\|\nabla I_{f}-max\{\nabla I_{ir},\nabla I_{vi}\}\|_{1},$ (18)
@@
where $\nabla$ represents the Sobel operator, which is used to calculate the gradient, $\|\cdot\|_{1}$ represents the $\ell_{1}$-norm, $H$ and $W$ denote the height and the width of the image, respectively, and $max\{\cdot,\cdot\}$ indicates the element-wise maximum operator.
@@
It is known that the maximum value selection-based pixel loss weakens the importance of pixels in one of the source images, while the average pixel loss may reduce the saliency of fusion results. Therefore, to achieve a better compromise between the preservation of important information and the enhancement of salient information, we adopt a segmented pixel loss function to train the proposed framework. In this work, the importance of a pixel, $\hat{p}$, is defined as the product of the pixel value and its gradient value:
$\hat{p}_{i,j}=\nabla I_{i,j}\cdot I_{i,j},$ (19)
@@
where $I_{i,j}$ represents the pixel value of the image $I$ at the $(i,j)$ position, while $\nabla I_{i,j}$ denotes the gradient value at the corresponding position. In this context, $\hat{P}$ and $\hat{Q}$ represent the pixel importance matrices of the infrared and visible light images, respectively.
@@
The importance of each pixel in the two source images is assessed and categorized into two segments as follows:
$\hat{P}_{top}=\{(i,j)\mid\hat{P}_{i,j}\geq\hat{p}_{ir}^{\alpha}\ \text{or}\ \hat{Q}_{i,j}\geq\hat{q}_{vi}^{\alpha}\}$ (20)
$\hat{P}_{res}=\{(i,j)\mid\hat{P}_{i,j}<\hat{p}_{ir}^{\alpha}\ \text{and}\ \hat{Q}_{i,j}<\hat{q}_{vi}^{\alpha}\}$
@@
where $\hat{p}_{ir}^{\alpha}$ and $\hat{q}_{vi}^{\alpha}$ represent the top $\alpha\%$ important pixels, and $\hat{P}_{top}$ denotes the first segment that consists of the most crucial pixels covering the top $\alpha\%$ important pixels from each source image. The remaining pixels are represented as the second segment, $\hat{P}_{res}$.
@@
As depicted above, we apply different pixel loss functions to each segment to meet the trade-off among pixels of different importance. The maximum value selection-based pixel loss is employed in $\hat{P}_{top}$, which can emphasize the saliency of these pixels in the fused image. Meanwhile, for $\hat{P}_{res}$, we use the pixel average loss to force the fused image to approximate each source image. The segmented pixel loss can be formulated as follows:
$\mathcal{L}_{top}=$ $\frac{1}{N_{top}}\sum_{(i,j)\in\hat{P}_{top}}\|I_{f}(i,j)-max(I_{ir}(i,j),I_{vi}(i,j))\|,$ (21)
$\mathcal{L}_{res}=$ $\frac{1}{2N_{res}}\sum_{(i,j)\in\hat{P}_{res}}\left(\|I_{f}(i,j)-I_{ir}(i,j)\|\right)$
$+\frac{1}{2N_{res}}\sum_{(i,j)\in\hat{P}_{res}}\left(\|I_{f}(i,j)-I_{vi}(i,j)\|\right)$
where $\mathcal{L}_{top}$ and $\mathcal{L}_{res}$ are the loss functions associated with the top $\alpha\%$ most important pixels and the remaining pixels, respectively. The number of pixels in $\hat{P}_{top}$, denoted as $N_{top}$, is given by $N_{top}=H\times W\times\alpha\%$, and the number of pixels in $\hat{P}_{res}$, denoted as $N_{res}$, is $N_{res}=H\times W\times(1-\alpha\%)$. The final pixel-wise loss function, $\mathcal{L}_{pixel}$, is defined as:
$\mathcal{L}_{pixel}=\mathcal{L}_{top}+\mathcal{L}_{res}.$ (22)
4 Experiments
@@
In this section, we first introduce the datasets and implementation settings. Then, the comparative experimental methods and objective evaluation metrics are described. After that, the experiment results and discussion are reported. Furthermore, we present the ablation experiments and the analysis of some hyperparameters and computational efficiency.
4.1 Datasets and Implementation Settings
In this study, we use three public datasets for our experiments: MSRS [53], RoadScene [29], and RGB-NIR Scene [54], see Tab. 1. The raw MSRS dataset contains 1,444 pairs of aligned IV images (training: 1083; test: 361). We randomly split the training dataset into 900 pairs for training and 180 pairs for validation. The training and validation image pairs are then randomly cropped in sample patches with a size of $128\times 128$. All test pairs in the MSRS dataset are used to verify the superior performance of the proposed method. To demonstrate the generalization ability, we chose 20 image pairs from the RoadScene dataset and 20 typical image pairs from the RGB-NIR Scene dataset.
Table 1: Datasets used for experiments (unit: image pairs)
Dataset Training Validation Test
MSRS 900 180 361
RoadScene 20
RGB-NIR Scene 20
To train the proposed ATFusion, we use the AdamW optimizer to update network parameters with an initial learning rate of $1\times 10^{-4}$. The learning rate drops to half of the previous one at 50, 100, 200, and 400 epochs. The batch size is set to 16. The parameters $\alpha$ and $\gamma$ are set to 20 and 1.0, respectively. Data augmentation techniques, including random horizontal flipping, noise injection, random cropping, and brightness adjustment, are applied to the original images with a probability of $20\%$. Additionally, a weight regularization constraint is incorporated to improve the model’s generalization and stability. The configurations for all experiments are computed with an NVIDIA GeForce RTX 4090 GPU and 64 GB RAM. The framework is programmed in PyTorch.
4.2 Comparative Methods and Evaluation Metrics
@@
Comparative Methods: To assess the superiority of our method, we select eight state-of-the-art image fusion methods for comparison, including DenseFuse [25], U2Fusion [29], RFN-Nest [27], SwinFuse [49], SwinFusion [19], LRRNet [55], DATFuse [18], and AEFusion [56]. The source codes of all compared methods are publicly available and all parameters are configured as suggested in the original papers. For a fair comparison, these state-of-the-art methods are trained on the same dataset to obtain fusion models, and the fusion results are generated on the same test datasets.
@@
Evaluation Metrics: Eight objective evaluation metrics are chosen to quantitatively evaluate the performance of all methods: average gradient (AG), entropy (EN) [57], feature mutual information (FMI) [58], standard deviation (SD) [59], spatial frequency (SF) [60], structural similarity (SSIM) [61], visual information fidelity (VIFF) [62], and $Q_{abf}$ [63]. AG measures the sharpness or clarity of an image. EN is an objective measure of the amount of information contained in an image. FMI evaluates the total of feature information transferred to the fused image from the source images. $Q_{abf}$ is a full-reference quality evaluation index estimating the degree of salient information from source images represented in the fused image. SD is a statistical theory-based standard deviation that reflects the degree of change in pixel brightness. SF is a measure based on gradient distribution, which represents the rate of change of the grayscale of the fused image. SSIM measures the closeness between the fused image and the source images from the brightness, structure, and contrast aspects. VIFF measures the quality of the fusion image according to the visual information fidelity. Higher values of these metrics correspond to better quality of the fused image.
4.3 Results and Discussion
Figure 5: Qualitative assessment of the compared VIF methods on the MSRS dataset.